OpenStudy (anonymous):

Alex made a sketch for a poster contest. The sketch measures 4 in. × 5 in. The scale factor from the sketch to the finished poster was 2 in. to 5 in. What were the dimensions of the poster?

OpenStudy (anonymous):

\[\frac{ 4 }{ 5 } \times \frac{ 5 }{ 5 } \div \frac{ 2 }{ 2 } = \frac{ 10 }{ 12.5 }\]

OpenStudy (anonymous):

Jane has a 5 × 7-inch photograph that she wants enlarged to a 10 × 14-inch photograph. What is the scale factor from the original photograph to the enlarged photograph?

OpenStudy (anonymous):

1 to 2, since everything is doubled.